brandy dawson and jason dodge start a motorcycle trip at the same point. both are traveling to san antonio a distance of about 400 kilometers. brandy rides 30 kilometers per hour faster than jason does. when brandy reaches her destination, jason has only traveled to austin, texas a distance of about 250 kilometers. determine the approx. speed of each motorcycle?
identify the variables, write the equation and solve algebraically

Answer by jorel555(1290)

Let r=jason's rate.
400/r+30=250/r
250(r+30)=400r
250r+7500=400r
150r=7500
r=50
r+30=80
Brandy is going 80 km/h, jason at 50 km/h!!!!!!!!!!!!!!